# Avenida Diagonal Las Torres 2640, 7941169, Peñalolén, Chile

### v1

@@ 5 Fully covariant counterterms in EGB AdS gravity in 6D
In even-dimensional Einstein-AdS gravity, bulk counterterms are given by the addition of the Euler density with a fixed coupling constant, which renders the total action finite for asymptotically conformally flat manifolds Olea:2005gb; Miskovic:2009bm; Anastasiou:2018rla; Anastasiou:2018mfk; Anastasiou:2020zwc. Nevertheless, additional divergences will appear if the asymptotic boundary is not Weyl-flat. A different renormalization scheme was later introduced, which complements the previous prescription and accounts for generic AAdS spacetimes Anastasiou:2020mik; Anastasiou:2022ljq; Anastasiou:2022wjq; Anastasiou:2023oro.
Interestingly, the use of topological densities of the Euler class as counterterms is not restricted only to Einstein-AdS gravity. In particular, it was shown that the same prescription turns EGB-AdS Kofinas:2006hr and, in general, Lovelock-AdS 0708.0782; Kofinas:2008ub gravity free from IR divergences for asymptotically conformally flat spaces. Nevertheless, its extension for generic AAdS spacetimes is still lacking. This section is intended to shed some light on this problem.
@@ 6.1 Boundary topology
Replacing the line element (82) in the Einstein-Gauss-Bonnet field equations (6), one finds that they are solved by the metric function Boulware:1985wk
@@
where $m$ is an integration constant associated to the mass, $\Sigma_{(4)}$ is the volume of the transverse section, while $k$ and $\sigma$ are related to its topology. Their specific values are summarized in Table (85).
@@
If the topology of the horizon is $\mathbb{S}^{4}$, this solution describes the one first found by Boulware and Deser in Ref. Boulware:1985wk. Thus, the solution in Eq. (84) will be referred to as Boulware-Deser topological black holes.
@@ 6.2 Asymptotic form
where $\ell_{\rm eff}\to\ell$ as $\alpha\to 0$, with $\ell_{\rm eff}$ being defined in terms of the cosmological constant via Eq. (9). Thus, in this limit, it is clear that this solution becomes the Schwarzschild-Tangherlini black hole Tangherlini:1963bw. On the other hand, the solution $f_{+}(r)$ is nonperturbative in $\alpha$. Henceforth, the analysis below focuses on the branch (87) for the sake of simplicity, denoted by $f_{-}(r)=f(r)$ from hereon. The nonperturbative case can be treated similarly. The asymptotic behavior of the metric function is
@@ 7 Conserved charges and thermodynamics
In order to compute the renormalized conserved charges for the topological Boulware-Deser-AdS black holes, the Noether-Wald formalism Wald:1993nt; Iyer:1994ys will be used. To this end, the Einstein-Gauss-Bonnet theory augmented by the fully covariant counterterms of Sec. 5 will be considered.
@@ 7.1 Noether-Wald charges
The Noether-Wald formalism uses the fact that the current associated with diffeomorphism invariance generated by a Killing vector field $\xi=\xi^{\mu}\partial_{\mu}$ is conserved on shell. The Noether current can be written locally as an exact form, whose prepotential can be expressed by Wald:1993nt; Iyer:1994ys
@@
Then, the asymptotic Noether-Wald charges associated with a Killing vector field $\xi$ can be obtained from Wald:1993nt; Iyer:1994ys
@@ 7.2 Quasilocal charges
Then, the quasilocal charges associated with the Killing vector $\xi=\xi^{\mu}\partial_{\mu}$ are defined as Balasubramanian:1999re
@@ 7.3 Black hole Thermodynamics
respectively. As is well-known, black hole entropy in Einstein-Gauss-Bonnet gravity does not follow the standard area law because it receives contributions from the higher-curvature terms Jacobson:1993xs; Cai:2001dz. Additionally, $S_{0}$ is a constant contribution to the entropy that does not affect the first law of thermodynamics. The latter has a topological origin, as it is related to the Euler characteristic of the black hole’s horizon via
@@ 8 Discussion and conclusions
In this work, Holographic Renormalization is performed for Einstein-Gauss-Bonnet AdS gravity in arbitrary dimensions, up to order $z^{5}$ in the Poincaré coordinate, for a generic boundary metric $g_{(0)ij}$. In particular, this procedure deals with divergences in AAdS spaces induced by a non-conformally flat boundary metric. On purpose, the corresponding counterterms are expressed in terms of the Weyl, Schouten and Cotton tensors, quantities of interest in the context of Conformal Calculus. They adequately account for nontrivial conformal properties of the boundary. This result contrasts with existing literature Brihaye:2008xu where, assuming covariance at the boundary, the same counterterms series as for Einstein gravity was used. The coefficients of this series were then fixed by demanding finiteness of the action for ACF black holes, which is not the case here. It is worthwhile to emphasize that the coefficients of the local counterterms in Eq. (44) are nonlinear in $\alpha$. When linearized in this parameter, the expressions match the ones obtained perturbatively by the Hamilton-Jacobi method in Ref. Liu:2008zf.
@@
Topological black holes with boundaries with a nontrivial conformal structure also serve to explore modifications of renormalization methods in higher-curvature AAdS gravity. As these conformal properties are reflected in divergent terms in the action in six and higher dimensions, in the case of six-dimensional EGB-AdS, the action is renormalized by fully covariant counterterms (see Section 5). Said counterterms extend the topological renormalization result of Ref. Kofinas:2006hr, by considering an additional term, $\Box\lvert W\rvert^{2}$, which cancels a divergence that appears in generic AAdS manifolds.
@@
When rewriting the bulk action in terms of a polynomial of the AdS curvature, it is easy to understand the role of $\Box\lvert W\rvert^{2}$. Basically, as shown in Eq. (73), the polynomial part of the renormalized action considers a quadratic and a cubic term in $F$. However, the cubic term is finite on its own. Therefore, the role of the extra counterterm is simply to cancel the divergence in the quadratic part of the action. This behavior mimics the situation for Einstein-AdS gravity when considering the Conformal Renormalization procedure, as shown in Ref. Anastasiou:2023oro. The reason behind this is the universality of next-to-leading order coefficient in the FG expansion Imbimbo:1999bj. Indeed, $g_{(2)ij}$ is fixed in terms of the Schouten tensor of $g_{(0)ij}$, what is a generic property of any AAdS spacetime, that can be shown even at a kinematic level. Particular features of EGB AdS appear at order $z^{4}$, where $g_{(4)ij}$ portrays an additional $\mathcal{W}_{(0)}^{2}$ contribution proportional to the GB coupling. However, such deformation is not enough to modify the conformal invariant structure in Eq. (4) until holographic order. As a consequence, one can borrow the form of the total derivative, which renders the action finite.
@@
As an application of this fully covariant renormalized action, Noether-Wald charges are derived, and explicitly computed for topological Boulware-Deser black holes. In doing so, the black hole mass, understood as internal energy, gets combined with the entropy and Helmholtz free energy to form the Quantum Statistical Relation (99) (see Gibbons:2004ai). Regarding quasilocal methods, the renormalized stress tensor in Eq. (47) also gives the correct conserved quantities for the AAdS solutions mentioned above. The use of boundary counterterms also reproduces the correct black hole thermodynamics. The holographic energy-momentum tensor for this family of solutions is traceless, as there is no conformal anomaly in odd-dimensional boundaries. Indeed, its perturbations describe an almost-conformal perfect fluid, where transport coefficients can be obtained in non-conformally flat backgrounds Kovtun:2004de.
@@
Renormalized gravity actions which are fully covariant in the bulk have a number of useful advantages, which have been explored to a certain extent. For instance, as shown in Ref. Anastasiou:2021jcv, the computation of the universal part of the holographic entanglement entropy for higher-curvature gravity theories Dong:2013qoa, such as Lovelock, simplifies considerably when the renormalized action is written in terms of spacetime-covariant objects. The implementation of $\Box\lvert W\rvert^{2}$ in codimension-2 would allow to obtain corrections to the Willmore energy or of the reduced Hawking mass in Einstein-AdS gravity Anastasiou:2024rxe. In addition to applications in holographic quantum information, in the context of the fluid/gravity correspondence, the covariant form of the renormalized EGB action may simplify the calculation of physical quantities in different coordinate frames Rangamani:2009xk; Bhattacharyya:2007vjd.
@@
Both quasilocal and Noether-Wald charges are intended to amend another notion of conserved quantities in AAdS spacetimes, which is Conformal Mass AshtekarMagnon1984; Ashtekar:1999jx. In the case of EGB-AdS gravity, the Conformal Mass definition, is given by the surface integral Pang:2011cs; Jatkar:2015ffa
@@
in terms of the normal vector to the boundary. This construction assumes that these components of the Weyl tensor are of holographic order, and in fact, this is the key argument to truncate Kounterterm charges in Einstein AdS Olea:2005gb; Olea:2006vd and in Einstein-Gauss-Bonnet AdS gravity Kofinas:2006hr to the linear order in the AdS curvature Jatkar:2014npa; Jatkar:2015ffa. However, note that the FG expansion of the electric part of the Weyl tensor
@@
Along this line, AMD-like formulas were derived for Lovelock AdS in Refs. Arenas-Henriquez:2017xnr; Arenas-Henriquez:2019rph. They are proportional to the degeneracy condition
@@
given in terms of the $\alpha_{p}$ couplings (each one corresponding to the Lovelock scalar of order $p$), and the maximal number of curvatures $N$. Due to the fact any theory of the Lovelock type is asymptotically Einstein (with a modified Newton’s constant), it is clear that the electric part of the Weyl tensor will turn Conformal Mass divergent when evaluated in non-ACF spacetimes Dotti:2005rc; Ray:2015ava. Dealing with this sort of infrared divergences will be the topic of future investigation. However, one may anticipate that, in even spacetime dimensions, additional boundary terms prescribed by higher-dimensional Conformal Invariants may play an essential role in producing renormalized action and charges. In doing so, one may conclude that the holographic stress tensor for an arbitrary Lovelock-AdS gravity is given by

### v2

@@ 5 Fully covariant counterterms in EGB AdS gravity in 6D
In even-dimensional Einstein-AdS gravity, bulk counterterms are given by the addition of the Euler density with a fixed coupling constant, which renders the total action finite for asymptotically conformally flat manifolds Olea (2005); Miskovic and Olea (2009); Anastasiou et al. (2018c, a, 2020). Nevertheless, additional divergences will appear if the asymptotic boundary is not Weyl-flat. A different renormalization scheme was later introduced, which complements the previous prescription and accounts for generic AAdS spacetimes Anastasiou et al. (2021c, 2022, 2023a, 2023b).
Interestingly, the use of topological densities of the Euler class as counterterms is not restricted only to Einstein-AdS gravity. In particular, it was shown that the same prescription turns EGB-AdS Kofinas and Olea (2006) and, in general, Lovelock-AdS Kofinas and Olea (2007, 2008) gravity free from IR divergences for asymptotically conformally flat spaces. Nevertheless, its extension for generic AAdS spacetimes is still lacking. This section is intended to shed some light on this problem.
@@ 6.1 Boundary topology
Replacing the line element (82) in the Einstein-Gauss-Bonnet field equations (6), one finds that they are solved by the metric function Boulware and Deser (1985)
@@
where $m$ is an integration constant associated to the mass, $\Sigma_{(4)}$ is the volume of the transverse section, while $k$ and $\sigma$ are related to its topology. Their specific values are summarized in Table 1. If the topology of the horizon is $\mathbb{S}^{4}$, this solution describes the one first found by Boulware and Deser in Ref. Boulware and Deser (1985). Thus, the solution in Eq. (84) will be referred to as Boulware-Deser topological black holes.
@@
Table 1: Properties of transverse sections.
@@ 6.2 Asymptotic form
where $\ell_{\rm eff}\to\ell$ as $\alpha\to 0$, with $\ell_{\rm eff}$ being defined in terms of the cosmological constant via Eq. (9). Thus, in this limit, it is clear that this solution becomes the Schwarzschild-Tangherlini black hole Tangherlini (1963). On the other hand, the solution $f_{+}(r)$ is nonperturbative in $\alpha$. Henceforth, the analysis below focuses on the branch (86) for the sake of simplicity, denoted by $f_{-}(r)=f(r)$ from hereon. The nonperturbative case can be treated similarly. The asymptotic behavior of the metric function is
@@ 7 Conserved charges and thermodynamics
In order to compute the renormalized conserved charges for the topological Boulware-Deser-AdS black holes, the Noether-Wald formalism Wald (1993); Iyer and Wald (1994) will be used. To this end, the Einstein-Gauss-Bonnet theory augmented by the fully covariant counterterms of Sec. 5 will be considered.
@@ 7.1 Noether-Wald charges
The Noether-Wald formalism uses the fact that the current associated with diffeomorphism invariance generated by a Killing vector field $\xi=\xi^{\mu}\partial_{\mu}$ is conserved on shell. The Noether current can be written locally as an exact form, whose prepotential can be expressed by Wald (1993); Iyer and Wald (1994)
@@
Then, the asymptotic Noether-Wald charges associated with a Killing vector field $\xi$ can be obtained from Wald (1993); Iyer and Wald (1994)
@@ 7.2 Quasilocal charges
Then, the quasilocal charges associated with the Killing vector $\xi=\xi^{\mu}\partial_{\mu}$ are defined as Balasubramanian and Kraus (1999)
@@ 7.3 Black hole Thermodynamics
respectively. As is well-known, black hole entropy in Einstein-Gauss-Bonnet gravity does not follow the standard area law because it receives contributions from the higher-curvature terms Jacobson and Myers (1993); Cai (2002). Additionally, $S_{0}$ is a constant contribution to the entropy that does not affect the first law of thermodynamics. The latter has a topological origin, as it is related to the Euler characteristic of the black hole’s horizon via
@@ 8 Discussion and conclusions
In this work, Holographic Renormalization is performed for Einstein-Gauss-Bonnet AdS gravity in arbitrary dimensions, up to order $z^{5}$ in the Poincaré coordinate, for a generic boundary metric $g_{(0)ij}$. In particular, this procedure deals with divergences in AAdS spaces induced by a non-conformally flat boundary metric. On purpose, the corresponding counterterms are expressed in terms of the Weyl, Schouten and Cotton tensors, quantities of interest in the context of Conformal Calculus. They adequately account for nontrivial conformal properties of the boundary. This result contrasts with existing literature Brihaye and Radu (2008) where, assuming covariance at the boundary, the same counterterms series as for Einstein gravity was used. The coefficients of this series were then fixed by demanding finiteness of the action for ACF black holes, which is not the case here. It is worthwhile to emphasize that the coefficients of the local counterterms in Eq. (44) are nonlinear in $\alpha$. When linearized in this parameter, the expressions match the ones obtained perturbatively by the Hamilton-Jacobi method in Ref. Liu and Sabra (2010).
@@
Topological black holes with boundaries with a nontrivial conformal structure also serve to explore modifications of renormalization methods in higher-curvature AAdS gravity. As these conformal properties are reflected in divergent terms in the action in six and higher dimensions, in the case of six-dimensional EGB-AdS, the action is renormalized by fully covariant counterterms (see Section 5). Said counterterms extend the topological renormalization result of Ref. Kofinas and Olea (2006), by considering an additional term, $\Box\lvert W\rvert^{2}$, which cancels a divergence that appears in generic AAdS manifolds.
@@
When rewriting the bulk action in terms of a polynomial of the AdS curvature, it is easy to understand the role of $\Box\lvert W\rvert^{2}$. Basically, as shown in Eq. (73), the polynomial part of the renormalized action considers a quadratic and a cubic term in $F$. However, the cubic term is finite on its own. Therefore, the role of the extra counterterm is simply to cancel the divergence in the quadratic part of the action. This behavior mimics the situation for Einstein-AdS gravity when considering the Conformal Renormalization procedure, as shown in Ref. Anastasiou et al. (2023b). The reason behind this is the universality of next-to-leading order coefficient in the FG expansion Imbimbo et al. (2000). Indeed, $g_{(2)ij}$ is fixed in terms of the Schouten tensor of $g_{(0)ij}$, what is a generic property of any AAdS spacetime, that can be shown even at a kinematic level. Particular features of EGB AdS appear at order $z^{4}$, where $g_{(4)ij}$ portrays an additional $\mathcal{W}_{(0)}^{2}$ contribution proportional to the GB coupling. However, such deformation is not enough to modify the conformal invariant structure in Eq. (4) until holographic order. As a consequence, one can borrow the form of the total derivative, which renders the action finite.
@@
As an application of this fully covariant renormalized action, Noether-Wald charges are derived, and explicitly computed for topological Boulware-Deser black holes. In doing so, the black hole mass, understood as internal energy, gets combined with the entropy and Helmholtz free energy to form the Quantum Statistical Relation (98) (see Gibbons et al. (2005)). Regarding quasilocal methods, the renormalized stress tensor in Eq. (47) also gives the correct conserved quantities for the AAdS solutions mentioned above. The use of boundary counterterms also reproduces the correct black hole thermodynamics. The holographic energy-momentum tensor for this family of solutions is traceless, as there is no conformal anomaly in odd-dimensional boundaries. Indeed, its perturbations describe an almost-conformal perfect fluid, where transport coefficients can be obtained in non-conformally flat backgrounds Kovtun et al. (2005).
@@
Renormalized gravity actions which are fully covariant in the bulk have a number of useful advantages, which have been explored to a certain extent. For instance, as shown in Ref. Anastasiou et al. (2021a), the computation of the universal part of the holographic entanglement entropy for higher-curvature gravity theories Dong (2014), such as Lovelock, simplifies considerably when the renormalized action is written in terms of spacetime-covariant objects. The implementation of $\Box\lvert W\rvert^{2}$ in codimension-2 would allow to obtain corrections to the Willmore energy or of the reduced Hawking mass in Einstein-AdS gravity Anastasiou et al. (2025). In addition to applications in holographic quantum information, in the context of the fluid/gravity correspondence, the covariant form of the renormalized EGB action may simplify the calculation of physical quantities in different coordinate frames Rangamani (2009); Bhattacharyya et al. (2008).
@@
Both quasilocal and Noether-Wald charges are intended to amend another notion of conserved quantities in AAdS spacetimes, which is Conformal Mass Ashtekar and Magnon (1984); Ashtekar and Das (2000). In the case of EGB-AdS gravity, the Conformal Mass definition, is given by the surface integral Pang (2011); Jatkar et al. (2015)
@@
in terms of the normal vector to the boundary. This construction assumes that these components of the Weyl tensor are of holographic order, and in fact, this is the key argument to truncate Kounterterm charges in Einstein AdS Olea (2005, 2007) and in Einstein-Gauss-Bonnet AdS gravity Kofinas and Olea (2006) to the linear order in the AdS curvature Jatkar et al. (2014, 2015). However, note that the FG expansion of the electric part of the Weyl tensor
@@
Along this line, AMD-like formulas were derived for Lovelock AdS in Refs. Arenas-Henriquez et al. (2017, 2019). They are proportional to the degeneracy condition
@@
given in terms of the $\alpha_{p}$ couplings (each one corresponding to the Lovelock scalar of order $p$), and the maximal number of curvatures $N$. Due to the fact any theory of the Lovelock type is asymptotically Einstein (with a modified Newton’s constant), it is clear that the electric part of the Weyl tensor will turn Conformal Mass divergent when evaluated in non-ACF spacetimes Dotti and Gleiser (2005); Ray (2015). Dealing with this sort of infrared divergences will be the topic of future investigation. However, one may anticipate that, in even spacetime dimensions, additional boundary terms prescribed by higher-dimensional Conformal Invariants may play an essential role in producing renormalized action and charges. In doing so, one may conclude that the holographic stress tensor for an arbitrary Lovelock-AdS gravity is given by